# Orbital orders under magnetic fields in cubic PrIr2Zn20

### v3

@@ IV.2 Phase diagram
At lower $T$, the induced $\Gamma$ point moment $\bm{\phi}_{0}$ favors the configuration III2q shown in Figs. 8(c) and 8(f), where $\bm{\phi}_{1}=\tfrac{1}{\sqrt{2}}\phi_{1}(1,1)$ and $\bm{\phi}_{2}=\tfrac{1}{\sqrt{2}}\phi_{2}(-1,1)$ with $\phi_{1}=\phi_{2}$. This configuration is stabilized by the fifth-order coupling discussed in the last part of Sec. IV.1. See also the last term in Eq. (39). In Fig. 6(a), an intermediate phase II2q toward the phase III2q appears at low temperatures. The quadrupole configurations in the phase II2q are “rotated” from those in the phase I2q with keeping $\bm{\phi}_{1}\perp\bm{\phi}_{2}$ as shown in Fig. 8(b). For smaller $b=0.05$ [Fig. 6(b)], the phase I, II, and the double-$\bm{q}$ orders shrink, while the phase III expands. The overall phase diagram, however, is unchanged.
Figure 7: $T$$H$ phase diagram for $\bm{H}\parallel[110]$, $J_{1}=1.0$, and $c=0.5$. (a) $(b,J_{2},J_{3},J_{4})=(0.1,0.0,0.0,1.0)$, (b) $(0.05,0.0,0.0,1.0)$, and (c) $(0.03,0.5,0.3,0.4)$, which correspond to those in Fig. 7(a), 7(b), and 7(c), respectively. Color represents $\Delta\phi\equiv|\phi_{1}-\phi_{2}|$, similarly to Fig. 6
Figure 8: Schematic orthogonal double-$\bm{q}$ order parameter configurations for (a) I2q: $(\theta_{1},\theta_{2})=(0,\pi/2)$ with $\phi_{1}>\phi_{2}$, (b) II2q: $(\theta_{1},\theta_{2})=(\theta,\theta+\pi/2)$ with $\phi_{1}>\phi_{2}$, and (c) III2q: $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$ and $\phi_{1}=\phi_{2}$. Arrows indicate the order parameters $\bm{\phi}_{1}$ (L1), $\bm{\phi}_{2}$ (L2), and $\bm{\phi}_{0}$ $(\Gamma)$. Four small circles represent the quadrupole moments at the $A$ sublattice real-space positions $\bm{r}=(0,0,0),(1,0,0),(\tfrac{1}{2},\frac{1}{2},0)$, and $(\tfrac{1}{2},\frac{1}{2},1)$. Real space L point quadrupole configurations for (d) $(\theta_{1},\theta_{2})=(0,\pi/2)$ and $\phi_{1}=\sqrt{3}\phi_{2}$ in the phase I2q, (e) $(\theta_{1},\theta_{2})=(\pi/24,13\pi/24)$ and $\phi_{1}=1.2\phi_{2}$ in the phase II2q, and (f) $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$, and $\phi_{1}=\phi_{2}$ in the phase III2q. Note that only L point quadrupole components without those at the $\Gamma$ point are shown. For guide to eyes, some of the nearest-neighbor bonds are connected by white bonds. Circles are the $A$ sublattice atoms and their colors represent the sublattice (red, green, blue, white). For clarity, the $B$ sublattice atoms are not drawn. Colored bonds running [01¯$1]$ 1 ] directions connects the $A$ sublattice sites with the same quadrupole moments.
As an example of more general cases, we show the phase diagram for $(b,J_{2},J_{3},J_{4})=(0.03,0.5,0.3,0.4)$ in Fig. 6(c). For this parameter set, the second-order coefficient $a_{\rm X}$ of the X point fields almost vanish at $T=0$ as shown in Fig. 1(c), the renormalization effects due to the X point fields grows rapidly at low temperatures. This causes the strong stabilization of the phase III2q in Fig. 6(c), showing reentrant behavior as a function of $H$. Note, however, the Landau theory is not applicable to such low temperature region far away from $T\sim T_{c}$. There is a thin region of the phase II2q. The phase boundary between the phase II2q and II is of the second order. Similarly, that between the small region of phase I2q and III is also second order. We note that the boundaries indicated by the dashed lines are not symmetry breaking phase boundary since the symmetry of the three double-$\bm{q}$ phases are the same as is evident from the real space configurations in Figs. 8(d)–8(f). See also the discussion in Appendix C.
@@
For $\bm{H}\parallel[110]$, a phase I2q appears at low-field region as shown in Fig. 7. In contrast to the phase I2q for $\bm{H}\parallel[001]$, the quadrupole configuration is $\bm{\phi}_{1}=(0,\phi_{1})$ and $\bm{\phi}_{2}=(\phi_{2},0)$ with $\phi_{1}>\phi_{2}$. This is natural since the neighboring single-$\bm{q}$ phase is the phase IV. The sharp increase of the critical field between I2q and IV phases in Fig. 7(c) is due to the fact that $a_{\rm X}$ is close to zero.
@@ V.1 Comparison with the experimental data of PrIr2Zn20
PrIr2Zn20 shows several phase transitions under magnetic field at low temperatures as reported in the early study.Onimaru et al. (2011) However, the order parameters have not been identified except for the $x^{2}-y^{2}$ type quadrupole order with the ordering vector at the L point for $\bm{H}\parallel[110]$ by the neutron scattering experimentIwasa et al. (2017). The analysis of this study is based on this experimental fact. Namely, we have assumed that the leading order parameter is the quadrupole moment at the L points. Constructing the symmetry-allowed free energy consisting of them and the related degrees of freedom, we have analyzed the phase diagram for $\bm{H}\parallel[001]$ in addition to that for $\bm{H}\parallel[110]$. Our analysis demonstrates that a $2z^{2}-x^{2}-y^{2}$ configuration with the ordering vector at the L point is stable at high-$T$ and up to intermediate magnetic field strength for $\bm{H}\parallel[001]$. This suggests that the $2z^{2}-x^{2}-y^{2}$ order can be a promising candidate for the high-$T$ phase found for $\bm{H}\parallel[001]$ in the experimental worksOnimaru et al. (2011); Kittaka et al. (2024). As for the order parameter at zero field, the single-$\bm{q}$ configuration is not stable while the double-$\bm{q}$ I2q is realized with $\phi_{1}=\phi_{2}$. For finite ${\bm{H}}\parallel[001]$, this double-$\bm{q}$ configuration is modified and $\bm{\phi}_{1}=(\phi_{1},0)$ and $\bm{\phi}_{2}=(0,\phi_{2})$ with $\phi_{1}>\phi_{2}$ is favored. $\phi_{2}$ increases as lowering $T$ and there appear phase II2q and III2q by simultaneously rotating $\bm{\phi}_{1}$ and $\bm{\phi}_{2}$. As the cubic anisotropy $b$ increases, the influence of the renormalized fourth-order coefficients becomes larger and favors the phase II2q and III2q.
@@
As magnetic field increases, the double-$\bm{q}$ order is destabilized and single-$\bm{q}$ orders take place as discussed in Sec. III. Thus, as increasing $\bm{H}\parallel[001]$, the double-$\bm{q}$, the single-$\bm{q}$ phase II with both $u_{\rm L}$ and $v_{\rm L}$ finite, and then the phase III with only $v_{\rm L}$ are realized. The trace of the phase boundaries between them have not been clearly observed in the experiments. However, there are several anomalies in the ultrasonic experimentsIshii et al. (2011). We hope the relation between our phase diagram and these anomalies can be clarified in future experimental studies. In addition to this, there is ambiguity about the stability of the phase III, whether the phase III extends toward the higher field as shown in Fig. 3(c) or occupies only small area [Fig. 3(a)]. This can be controlled by the cubic coupling $b$. For $\bm{H}\parallel[110]$, there is no high-$T$ phase in contrast to the case for $\bm{H}\parallel[001]$ in the experiments.Onimaru et al. (2011); Kittaka et al. (2024) This point is consistent with our results. Other anomalies observed in the ultrasound experimentsIshii et al. (2011) similar to those for $\bm{H}\parallel[001]$ are also interesting to examine in more detail in the future works. Since the anomalies are tiny and thus the comparison between the experiments and the present theory is not straightforward, while there is a phase boundary between the phase I2q and IV in Fig. 7.
@@
We also note that the phase diagram as a function of magnetic field from $\bm{H}\parallel[001]$ to $\bm{H}\parallel[110]$ is described in a function of angle $\Theta$, where $\bm{H}=\tfrac{1}{2}H(\sin\Theta,\sin\Theta,2\cos\Theta)$. In Eq. (12), the effective field to the quadrupole moments is written as $\tilde{\bm{h}}=gH^{2}(3\cos^{2}\Theta-1,0)$. Thus, the critical field of quadrupole orders for the angle $\Theta$, $H_{c}(\Theta)$ is expected to be
@@
Such behavior can be one of strong evidences for quadrupole orders. We hope the future experimental studies on the phase diagram at finite fields clarify this point.
V.2 Future perspectives
@@
We have shown that single- and double-$\bm{q}$ quadrupole orders are realized in our study as a phenomenological model for PrIr2Zn20. The orthogonal double-$\bm{q}$ orders are favored by the effective fourth-order terms arising from the renormalization due to the X point fields, while the quadruple ones are not. Although we are interested in the orbital orders at the L points in this paper, quadruple-$\bm{q}$ orders at incommensurate ordering vectors along [111] directions has been studied in cubic magnets, showing hedgehog type topological magnetic structure.Okumura et al. (2020, 2022) In principle, such exotic phases are also possible for the orbital model discussed in this paper when the exchange coupling favors incommensurate ordering vectors by introducing interaction frustration or anisotropic interaction ignored in this studyHattori et al. (2024). In this sense, it is interesting to explore possibilities of incommensurate orbital orders as the candidates for the high-$T$ and intermediate field phase instead of the phase I proposed in this paper as studied in magnetic ordersD. Bergman and Balents (2007); S. Gao and Rüegg (2017); Rosales et al. (2022).
@@
For quantitative comparison between the theoretical and experimental results, it is also needed to incorporate the actual crystal electric field states in the real materials. They vary under magnetic fields and thus the parameters in the free energy can change. For theoretical or computational aspects, it is also interesting to take into account the fluctuation effects by using e.g., classical Monte Carlo simulation employed in the previous studies on Pr-based materials.Hattori and Tsunetsugu (2016); Freyer et al. (2018, 2020) The thermal fluctuation may alter the mean-field phase diagram and to check their stability is important issue as one of our future problems.
@@ VI Summary
We have analyzed quadrupole orders with its modulation characterized by the wavenumber at the L points. Their Landau free energy is constructed under magnetic fields $\bm{H}$. We find that the phase diagrams show strong anisotropy. We have clarified that $x^{2}-y^{2}$ type quadrupole order is stabilized along $\bm{H}\parallel[110]$, which is consistent with that proposed in the neutron scattering experiment. We also propose that the unidentified high-temperature phase for $\bm{H}\parallel[001]$ in the experiments is a collinear $2z^{2}-x^{2}-y^{2}$ order and the orthogonal double-$\bm{q}$ configurations are stable at low fields. These results offers further experimental examination of the order parameters in PrIr2Zn20 and related compounds.
@@ Acknowledgment
The authors thank T. Onimaru T. Ishitobi, and S. Kittaka for fruitful discussion. This work was supported by JSPS KAKENHI (Grant Nos. JP23K20824 and JP23H04869).

### v4

@@ IV.2 Phase diagram
There is another double-$\bm{q}$ order labeled by II2q in Figs. 6(a) and 6(b). The quadrupole configurations in the phase II2q are “rotated” from those in the phase I2q with keeping $\bm{\phi}_{1}\perp\bm{\phi}_{2}$ as shown in Fig. 7(b). For smaller $b=0.05$ [Fig. 5(b)], the phase I, II, and the double-$\bm{q}$ orders shrink, while the phase III expands. The overall phase diagram, however, is unchanged.
@@
For $\bm{H}\parallel[110]$, the phase I2q appears at low-field region as shown in Figs. 6(c) and 6(d), while the phase II2q does not appear. In contrast to the phase I2q for $\bm{H}\parallel[001]$, the quadrupole configuration is $\bm{\phi}_{1}=(0,\phi_{1})$ and $\bm{\phi}_{2}=(\phi_{2},0)$ with $\phi_{1}>\phi_{2}$. This is natural since the neighboring single-$\bm{q}$ phase is the phase IV. The fact that $\phi_{1}$ is larger than $\phi_{2}$ explains why there is no II2q for $\bm{H}\parallel[110]$. The coupling $\lambda_{c}^{(2)}$ in Eq. (15) favors $\bm{\phi}_{n}\perp\bm{\phi}_{0}$. This suggests that the configuration in the phase I2q for $\bm{H}\parallel[110]$ is stable since the larger $\bm{\phi}_{1}$ is perpendicular to $\bm{\phi}_{0}$, while parallel in the phase I2q for $\bm{H}\parallel[110]$.
Figure 8: TTHH phase diagram for $J_{1}=1.0$, $c=0.5$ and $(b,J_{2},J_{3},J_{4})=(0.03,0.5,0.3,0.4)$ for (a) $\bm{H}\parallel[001]$ and (b) $\bm{H}\parallel[110]$. Color represents $\Delta\phi\equiv|\phi_{1}-\phi_{2}|$, similarly to Fig. 6. (c) Schematic orthogonal double-$\bm{q}$ order parameter configurations in the phase III2q for $(\theta_{1},\theta_{2})=(\pi/4,3\pi/4)$ with $\phi_{1}=\phi_{2}$ (a) in the uuvv space and (b) in the real space. Colored bonds represent the uniform alignment of $\bm{\phi}^{A}(\bm{r})$ similarly to those in Figs. 7(b) and 7(d).
As an example of more general cases, we show the phase diagram for $(b,J_{2},J_{3},J_{4})=(0.03,0.5,0.3,0.4)$ in Fig. 8. At low temperatures, the induced $\Gamma$-point moment $\bm{\phi}_{0}$ favors the configuration III2q shown in Figs. 8(c) and 8(d), where $\bm{\phi}_{1}=\tfrac{1}{\sqrt{2}}\phi_{1}(1,1)$ and $\bm{\phi}_{2}=\tfrac{1}{\sqrt{2}}\phi_{2}(-1,1)$ with $\phi_{1}=\phi_{2}$. This configuration is stabilized by the fifth-order coupling discussed in the last part of Sec. IV.1. See also the last term in Eq. (39).
For this parameter set, the second-order coefficient $a_{\rm X}$ of the X-point fields almost vanish at $T=0$ as shown in Fig. 2(d), the renormalization effects due to the X point fields grows rapidly at low temperatures. This causes the strong stabilization of the phase III2q in Fig. 8(a) for $\bm{H}\parallel[001]$, showing reentrant behavior as a function of $H$. For $\bm{H}\parallel[110]$, the sharp increase of the critical field between I2q and IV phases in Fig. 8(b) is also due to the fact that $a_{\rm X}$ is close to zero. Note that the Landau theory is not applicable to such low temperature region far away from $T\sim T_{c}$. There is a thin region of the phase II2q for $\bm{H}\parallel[001]$. The phase boundary between the phase II2q and II is of the second order. Similarly, that between the small region of phase I2q and III is also second order. We note that the boundaries indicated by the dashed lines are not symmetry breaking phase boundary since the symmetry of the three double-$\bm{q}$ phases are the same as is evident from the real space configurations in Figs. 7(b), 7(d), and 8(d). See also the discussion in Appendix C.
@@ V Discussions
So far, we have shown how the L point quadrupole orders evolve under magnetic fields by using the Landau analysis. In this section, we will compare these results with the experimental data of PrIr2Zn20 and the related compounds. We also discuss the similarity between other quadrupole orders in Pr-based 1-2-20 compounds such as the Néel type antiferroic orders and those at the X points from theoretical point of view. In the last part of this section, we will comment on the future perspectives toward full understanding of the quadrupole orders in PrIr2Zn20.
@@ V.1 Comparison with the experimental data of PrIr2Zn20
PrIr2Zn20 shows several phase transitions under magnetic field at low temperatures as reported in the early study [31]. However, the order parameters have not been identified except for the $x^{2}$-$y^{2}$ type quadrupole order with the ordering wave number at the L point for $\bm{H}\parallel[110]$ by the neutron scattering experiment [34]. The analysis of this study is based on this experimental fact. Namely, we have assumed that the leading order parameter is the quadrupole moment at the L points. Constructing the symmetry-allowed free energy consisting of them and the related degrees of freedom, we have analyzed the phase diagram for $\bm{H}\parallel[001]$ in addition to that for $\bm{H}\parallel[110]$. Our analysis demonstrates that a $2z^{2}$-$x^{2}$-$y^{2}$ configuration with the ordering wave number at the L point is stable at high-$T$ and up to intermediate magnetic field strength for $\bm{H}\parallel[001]$. This suggests that the $2z^{2}$-$x^{2}$-$y^{2}$ order can be a promising candidate for the high-$T$ phase found for $\bm{H}\parallel[001]$ in the experimental works [31, 35]. As for the order parameter at zero field, the single-$\bm{q}$ configuration is not stable while the double-$\bm{q}$ I2q is realized with $\phi_{1}=\phi_{2}$. For finite ${\bm{H}}\parallel[001]$, this double-$\bm{q}$ configuration is modified and $\bm{\phi}_{1}=(\phi_{1},0)$ and $\bm{\phi}_{2}=(0,\phi_{2})$ with $\phi_{1}>\phi_{2}$ is favored. $\phi_{2}$ increases as lowering $T$ and there appear phase II2q and III2q by simultaneously rotating $\bm{\phi}_{1}$ and $\bm{\phi}_{2}$. As the cubic anisotropy $b$ increases, the influence of the renormalized fourth-order coefficients becomes larger and favors the phase II2q and III2q.
@@
As magnetic field increases, the double-$\bm{q}$ order is destabilized and single-$\bm{q}$ orders take place as discussed in Sec. III. Thus, as increasing $\bm{H}\parallel[001]$, the double-$\bm{q}$, the single-$\bm{q}$ phase II with both $u_{\rm L}$ and $v_{\rm L}$ finite, and then the phase III with only $v_{\rm L}$ are realized. The trace of the phase boundaries between them has not been clearly observed in the experiments. However, there are several anomalies in the ultrasonic experiments [32]. We hope the relation between our phase diagram and these anomalies can be clarified in future experimental studies. In addition to this, there is ambiguity about the stability of the phase III, whether the phase III extends toward the higher field as shown in Fig. 4(c) or occupies only small area [Fig. 4(a)]. This can be controlled by the cubic coupling $b$. For $\bm{H}\parallel[110]$, there is no high-$T$ phase in contrast to the case for $\bm{H}\parallel[001]$ in the experiments [31, 35]. This point is consistent with our results. Other anomalies observed in the ultrasound experiments [32] similar to those for $\bm{H}\parallel[001]$ are also interesting to examine in more detail in the future works. Since the anomalies are tiny and thus the comparison between the experiments and the present theory is not straightforward, while there is a phase boundary between the phase I2q and IV in Figs. 6(c), 6(d), and 8(b), irrespectively to the detail of the parameters.
@@
We also note that the phase diagram as a function of magnetic field from $\bm{H}\parallel[001]$ to $\bm{H}\parallel[110]$ is described in a function of angle $\Theta$, where $\bm{H}=\tfrac{1}{\sqrt{2}}H(\sin\Theta,\sin\Theta,\sqrt{2}\cos\Theta)$. In Eq. (12), the effective field to the quadrupole moments is written as $\tilde{\bm{h}}=gH^{2}(3\cos^{2}\Theta-1,0)$. Thus, the critical field of quadrupole orders for the angle $\Theta$, $H_{c}(\Theta)$ is expected to be
@@
Such behavior can be one of strong evidences for quadrupole orders and indeed have been observed recently [41].
V.2 Comparison with quadrupole orders at the $\Gamma$ point and the X points
In the previous studies, quadrupole orders in Pr-based 1-2-20 compounds have been theoretically analyzed in several types of ordering wavenumbers. In the early stage, simple Néel type quadrupole orders [21, 23] are analyzed and TTHH phase diagrams show similar order parameter switching as discussed in the analyses of single-$\bm{q}$ sates shown in Fig. 4. This is because the form of the Landau’s free energy for both cases are almost the same when the multiple-$\bm{q}$ configurations are neglected.
For PrTi2Al20, the experiments clarify the order is ferroic [20, 24, 25] and its unusual behavior in the quadrupole switching under magnetic fields is analyzed by field-induced interactions unique to the quadrupole system [24]. In this study such induced interactions are not taken into account. We do not consider they are important in PrIr2Zn20 since they require specific electronic properties such as existence of small pocket Fermi surfaces.
For PrV2Al20, there are two transitions at zero magnetic field. The high-temperature transition is considered to be a quadrupole order, while ferroic octupole moments emerge below the lower transition temperature [42, 27]. The Landau analysis for a model including the octupole moment in addition to the quadrupole moments [28] and Monte Carlo simulations with extended interactions have been done [29, 26]. These theories explain the finite octupole moment as a consequence of symmetry breaking of ferroic octupole order. The lack of microscopic observation of both quadrupole and octupole moments (even their ordering wavenumbers) prevents theoretical investigation from further analyses. As an alternative approach, it has been pointed out that if the high-temperature transition is that into triple-$\bm{q}$ quadrupole order at the X points suggested by the band structure calculation [43], the lower temperature transition is expected to be triple-$\bm{q}$ octupole order at the same wavenumbers with induced ferroic octupole moments [30].
The physics of the L point quadrupole order is distinct from the ferroic and the X-point quadrupole moments in particular in the absence of cubic invariant consisting solely of the order parameters since the sum of three ordering wavenumbers lead to ${\bm{k}}_{1}+{\bm{k}}_{2}+{\bm{k}}_{3}\neq{\bm{0}}$. Although the Néel type order parameters do not possess such cubic invariant, they have trivially no multiple-$\bm{q}$ orders. In this sense, the present results including double-$\bm{q}$ quadrupole orders and the phase transitions from/to the single-$\bm{q}$ orders are unique properties for the L point quadrupole systems.
V.3 Future perspectives
@@
We have shown that single- and double-$\bm{q}$ quadrupole orders are realized in our study as a phenomenological model for PrIr2Zn20. The orthogonal double-$\bm{q}$ orders are favored by the effective fourth-order terms arising from the renormalization due to the X-point fields, while the quadruple ones are not. Although we are interested in the orders at the L points in this paper, quadruple-$\bm{q}$ orders at incommensurate ordering wave numbers along [111] directions has been studied in cubic magnets, showing hedgehog type topological magnetic structure [44, 45]. In principle, such exotic phases are also possible for the orbital model discussed in this paper when the exchange coupling favors incommensurate ordering wave numbers by introducing interaction frustration or anisotropic interaction ignored in this study [46]. In this sense, it is interesting to explore possibilities of incommensurate orders as the candidates for the high-$T$ and intermediate field phase instead of the phase I proposed in this paper as studied in magnetic orders [40, 47, 48].
@@
For quantitative comparison between the theoretical and experimental results, it is also needed to incorporate the actual crystal electric field states in the real materials. They vary under magnetic fields and thus the parameters in the free energy can change. Recent ultrasonic experiments [49] detected rotational effects and its microscopic analysis is also interesting. For theoretical or computational aspects, it is also interesting to take into account the fluctuation effects by using e.g., classical Monte Carlo simulation employed in the previous studies on Pr-based materials [23, 29, 26]. The thermal fluctuation may alter the mean-field phase diagram and to check their stability is important issue as one of our future problems [50]. For example, our analysis of mode-mode coupling is restricted to those with the X-point modes via the cubic terms. This may cause overestimation of double-$\bm{q}$ orders. In this sense, the numerical simulation can clarify the effects of these fluctuations. The influence of the non-Fermi liquid properties [20, 33, 51, 18] on the symmetry breaking, which is completely ignored in this study, is also nontrivial and important subject.
@@ VI Summary
We have analyzed quadrupole orders with its modulation characterized by the wavenumber at the L points. Their Landau free energy is constructed under magnetic fields $\bm{H}$. We find that the phase diagrams show strong anisotropy. We have clarified that $x^{2}$-$y^{2}$ type quadrupole order is stabilized along $\bm{H}\parallel[110]$, which is consistent with that proposed in the neutron scattering experiment. We also propose that the unidentified high-temperature phase for $\bm{H}\parallel[001]$ in the experiments is a collinear $2z^{2}$-$x^{2}$-$y^{2}$ order and the orthogonal double-$\bm{q}$ configurations are stable at low fields. These results offer further experimental examination of the order parameters in PrIr2Zn20 and related compounds.
@@ Acknowledgment
The authors thank T. Onimaru, T. Ishitobi, and S. Kittaka for fruitful discussion. This work was supported by JSPS KAKENHI (Grant No. JP23K20824, No. JP23H04866, and No. JP23H04869).
Figure 9: (a) Double-$\bm{q}$ configuration $\mathbb{E}_{2}$. Colored thick tubes connect the same $\mathbb{E}_{2}$’s. Thin bonds indicate the nearest-neighbor bonds of diamond structure. Arrows indicate the direction of ${\bm{k}}_{1}$ and ${\bm{k}}_{2}$. (b) Quadruple-$\bm{q}$ configuration $\mathbb{E}_{4}$, where only the $A$-sublattice sites are shown by spheres. (c) Symmetric quadruple-$\bm{q}$ configuration for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,0,0)$ and (d) for $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(0,0,\pi/2,3\pi/2)$. For (d), there are four kinds of ferroic chains. Two of them are similar to the double-$\bm{q}$ configuration for $x=n$ ($n:$ integers) planes with the $2z^{2}$-$x^{2}$-$y^{2}$ orbital, while the other two chains for $x=n+\tfrac{1}{2}$ plane with the $x^{2}$-$y^{2}$ orbital run along ${\bm{k}}_{1}+{\bm{k}}_{2}$ direction.